OpenStudy (anonymous):

f(x)= (x^2 + 1)^2 g(x)= (1)/(sqrt(x+5)) what is f of g what is g of f please show work for both thanks in advance

OpenStudy (anonymous):

(fog)= ( (1/sqrt.x+5)^2 +1)^2 = (x^2 +12x +36) / (x^2 +10x+ 25) I just inserted g(x) into x of f(x) (gof)= 1/(sqrt.((x^2 +1)^2 ) +5)

OpenStudy (anonymous):

for the second one, just multiply the numerator and deonominator by the conjugate of the squareroot